RhondaSommer (rhondasommer):

The percent composition of acetic acid, HC2H3O2, is: 20% carbon, 2% hydrogen, 27% oxygen 40% carbon, 7% hydrogen, 53% oxygen 23% carbon, 7% hydrogen, 70% oxygen Cannot be determined from the information given I have no idea, @sweetburger can you help?

OpenStudy (sweetburger):

So percent composition is basically composition by mass. The total mass of acetic acid is 60.05 grams. I will do Hydrogen for you There are 4 hydrogens in acetic acid so 4*(molar mass of Hydrogen) / 60.05 so it would be roughly 4/60 x 100% = 6.5% H which rounds up to 7% Hydrogen. Just repeat the process with O and C.

RhondaSommer (rhondasommer):

so how many carbons are in acetic acid?

RhondaSommer (rhondasommer):

@sweetburger

OpenStudy (sweetburger):

2

RhondaSommer (rhondasommer):

so 3 carbon?

OpenStudy (sweetburger):

no only 2 carbons

OpenStudy (sweetburger):

2 oxygen as well

RhondaSommer (rhondasommer):

i means like 3% of carbon

RhondaSommer (rhondasommer):

2/60 = 1/30 *100 = roughly 3%

OpenStudy (sweetburger):

noo its molar mass of 2 carbons divided by the total molar mass of aceitic acid so like 24/60

RhondaSommer (rhondasommer):

wait what

OpenStudy (sweetburger):

mass composition is the finding what percent by mass does a single type of atom make up a compound. For example NaCl the Na atom does not make up 50% of this compounds mass instead it makes up (mass of Na)/(mass of NaCl) = roughly 40% Na by mass

RhondaSommer (rhondasommer):

so carbon would be 24 (ish) then it would be 40%

OpenStudy (sweetburger):

yes exactly

RhondaSommer (rhondasommer):

so B would be the answer right?

OpenStudy (sweetburger):

Yea looks like it